If 15.0 liters of oxygen at STP are heated to 500 o C, what will be

the new volume of gas if the pressure is also increased to 1620

mmHg?


1
Expert's answer
2022-06-13T03:09:03-0400

n(O2) = V / 22,4 mol/l = 15 / 22,4 = 0,67 mol


using the Mendeleev clapeyron equation, we find the volume of the gas under the new conditions:


PV = nRT

Where,

P — pressure = 1620 mmHg = 216 kPa

V — volume (liters) = ?

n — amount of gas = 0,67 mol

T — temperature = 500°C = 773 K

R — universal gas constant = 8,31


"V = \frac {nRT}{P} = \frac {0,67*8,31*773}{216}= 19,9 ~l"


Answer: 19,9 liters